Imogene's car traveled 280 mi averaging a certain speed. If the car had gone 5 mph faster, the trip would have taken 1 hour less. Find the average speed.




i let r= rate of average speed
i let r+5= rate could have been 1 hour less
280/r + 280/(r+5)= 280/r + 280/(r+5)

Answer by mananth(16946)

Imogene's car traveled 280 mi averaging a certain speed. If the car had gone 5 mph faster, the trip would have taken 1 hour less. Find the average speed.
..
let rate be x
if rate is increase by 5 the new rate = x+5 mph
Distance = 280 miles
...
Time taken by old rate - time taken by new rate =1
...
280/x-280/(x+5)=1
280(x+5)-280x /x(x+5) =1
280x+1400-280x=x(x+5)
1400 = x^2+5x
x^2+5x-1400=0
x^2++40x-35x-1400=0
x(x+40)-35(x+40)=0
(x+40)(x-35)=0
So x = 35 mph the original rate